3x-6=2y
If x is 8, what point (ordered pair), satisfies the equation?

Answer by ewatrrr(24785)


Hi,
*Note: "ordered pair"(x,y) represents the x value and corresponding y value that satisfys the equation given.
3x-6=2y
(3/2)x - 3 = y
If x is 8, what point (ordered pair), satisfies the equation?
substituting x = 8 into the equation
(3/2)x - 3 = y
(3/2)8 - 3 =12-3 = 9
ordered pair (8,9) satisfys the equation